The Smith Trucking Company claims that the average weight of its delivery trucks when fully loaded is 6000 pounds with a standard deviation of 120 pounds. 36 trucks are selected at random and their weights recorded. Within what limits will the average weights of 90% of the 36 trucks lie?


1
Expert's answer
2022-04-25T17:56:54-0400

"CI=\mu \plusmn z\frac{\sigma}{\sqrt{n}}"

"Z=1.645" from z value table "\alpha= (1-0.9)/2"

"CI=6000 \plusmn 1.645 \frac{120}{\sqrt{36}}=6000 \plusmn 32.9"